OpenStudy (anonymous):

A third grade class had a spelling bee with 6-letter and 7-letter words. They used 30 words that had a total of 190 letters. Which system of linear equations represents the problem? A.7x+6y=30 x+y=190 B.7x+6y=30 x+y=30 C.6x+7y=160 x+y=30 D.6x+y=30 x+7y=190

OpenStudy (anonymous):

x = 6letter y = 7letter x + 7 = 30 6x + 7y = 190 D

OpenStudy (anonymous):

No, Not D

OpenStudy (anonymous):

C, but replace the 160 with 190, I guess that was your mistake.

OpenStudy (anonymous):

C.6x+7y=160 x+y=30